# An Empirical Study of Conformal Prediction in LLM with ASP Scaffolds for Robust Reasoning

### v1

@@ Abstract
In this paper, we examine the use of Conformal Language Modelling (CLM) alongside Answer Set Programming (ASP) to enhance the performance of standard open-weight LLMs on complex multi-step reasoning tasks. Using the StepGame dataset, which requires spatial reasoning, we apply CLM to generate sets of ASP programs from an LLM, providing statistical guarantees on the correctness of the outputs. Experimental results show that CLM significantly outperforms baseline models that use standard sampling methods, achieving substantial accuracy improvements across different levels of reasoning complexity. Additionally, the LLM-as-Judge metric enhances CLM’s performance, especially in assessing structurally and logically correct ASP outputs. However, calibrating CLM with diverse calibration sets did not improve generalizability for tasks requiring much longer reasoning steps, indicating limitations in handling more complex tasks.
@@ 1 Introduction
Building automated reasoning systems is one of the cornerstone goals of research in artificial intelligence, with wide-reaching applications in a variety of domains garnering interest from the research community. LLMs have demonstrated remarkable capabilities in various natural language processing tasks Min et al. (2021), yet exhibit critical limitations in a variety of reasoning-related tasks, particularly in multistep and spatially complex scenarios Yang et al. (2023); Li et al. (2024). Although these models process and generate natural language with unprecedented fluency, their reasoning capabilities are still limited by several key constraints.
@@
Most LLMs frequently struggle with verifiable, robust and interpretable reasoning. Experimental studies have shown significant variability in LLM responses to identical reasoning related questions which highlights inherent instability Mirzaee & Kordjamshidi (2023). For instance, sampling from an LLM’s responses to reasoning-based queries can produce widely varying answers with significantly divergent reasoning traces. This undermines their reliability in critical decision-making contexts Ji et al. (2023). Moreover, tasks that require multiple sequential reasoning steps, such as complex spatial reasoning, often expose fundamental weaknesses in LLM’s systematic reasoning capabilities Shi et al. (2022).
@@
Recent research has increasingly focused on neural-symbolic approaches that combine the strengths of LLMs in learning from data and rigorous reasoning capabilities of symbolic systems Mao et al. (2019); Yu et al. (2024). Specifically, Answer Set Programming (ASP), a type of declarative logic programming, has emerged as a promising framework for enhancing LLMs’ reasoning capabilities (Yang et al., 2023). Recent studies have explored integrating LLMs with ASP to transform natural language into logical representations, particularly for spatial reasoning tasks (Yang et al., 2023; Li et al., 2024).
@@
However, existing approaches face significant challenges in generating reliable intermediate scaffolds – the logical representations that bridge natural language inputs and systematic symbolic reasoning systems. The translation of natural language into formal logic programs introduces substantial uncertainty, as LLMs can produce inconsistent or semantically incorrect translations Li et al. (2024). This variability may become a significant weakness in this reasoning pipeline, potentially compromising the reliability of downstream reasoning tasks Mirzaee & Kordjamshidi (2023).
@@
In this paper, we present an approach that exploits the recently proposed Conformal Language Modelling (CLM) Quach et al. (2024) to systematically address these scaffold generation challenges. Conformal prediction (CP), the foundational framework for CLM, offers unique advantages in uncertainty quantification Vovk et al. (2005); Angelopoulos et al. (2022). CLM has been applied to natural language generation tasks like summarisation and question-answering Quach et al. (2024). Although recent advances in conformal prediction for language generation show promise, significant questions remain regarding the scalability, generalisability, and robustness of these approaches across reasoning domains Campos et al. (2024). In this paper, we present a thorough analysis of these open challenges through an empirical lens. We assess how calibration in CP affects generalisability on multi-hop reasoning tasks and evaluate the efficacy of adapting metrics for formal language generation (ASP). Additionally, we identify common error types in the ASP scaffolded LLM based reasoning pipelines and examine how these propagate. Our findings aim to enhance the reliability and performance of neuro-symbolic reasoning systems, contributing to the development of more robust and interpretable reasoning frameworks.
@@ 2 Background
We introduce the main components used in our work: conformal prediction, conformal language modelling and answer set programming.
@@
Conformal Prediction is a statistical framework introduced by Vovk et al. (2005) that provides a method for constructing prediction regions with finite-sample guarantees under minimal assumptions. CP can be described as a sampling method which, under certain mild assumptions, guarantees that samples taken from any distribution using this method will be ‘correct’ with a controlled probability. This theory has been well-received in machine learning, where providing guarantees on the correctness of predictions helps improve performance, especially since inference using neural networks involves sampling from theoretically unknown distributions Campos et al. (2024). CP has already proven its utility and efficacy in various domains, such as in computer vision tasks Hechtlinger et al. (2018) and drug discovery Cortés-Ciriano & Bender (2019).
@@
Conformal Language Modelling (CLM), developed in Quach et al. (2024), is a recent extension of CP allowing for conformal sampling of language models. CLM proposes a conformal approach to generating sets of responses with language models, allowing for statistical guarantees in the vast and complex output space of natural language generation. The main contribution of CLM is an algorithm that can control a specific set-valued risk – the expected indicator loss that the generated set contains at least one acceptable response according to a predefined admission function—while simultaneously maximizing other quantities of interest such as quality, diversity, and confidence. However, it remains unclear how CLM affects generalisability, particularly because it has been primarily tested on natural language generation tasks like summarisation, and report generation, where verifying the accuracy of models is not always straightforward Campos et al. (2024). Applying CLM to domains with rigorous notions of correctness, such as formal language generation, can provide better insights into its handling of generalisability and potential limitations.
@@
Answer set programming is a type of declarative programming which is commonly used as a formal representation of knowledge, constraint satisfaction tasks, and formal common-sense reasoning. Answer set programs consist of finite sets of facts and rules that are used to generate answer sets under stable model semantics, which allows for efficient encoding of complex problems, notably in domains like graph search and combinatorial optimization Lifschitz (2019). ASP’s utility extends to various industrial applications due to its expressive power and computational efficiency, for instance, in e-tourism Ricca et al. (2010), high-performance computing scheduling Gamblin et al. (2022), configuration management, diagnosis, and planning Erdem et al. (2016). Further, ASP has been used to bolster spatial reasoning in LLMs through neural-symbolic integration Yang et al. (2023).
@@ 3 Formal introduction to CLM
At its core, CLM functions as a calibrated rejection sampling algorithm, allowing one to form conformal sets of samples, $\mathcal{C}(X)$, from an LLM, given a prompt $X$, such that the expected loss of $\mathcal{C}(X)$ is guaranteed to be within a chosen bound. Specifically, for a given loss function $L$ and error tolerance $\delta\in(0,1)$, CLM ensures there exists a risk tolerance $\varepsilon\in(0,1)$ such that for any prompt $X$, the conformal set $\mathcal{C}(X)$ satisfies the Equation 1 with probability at least $1-\delta$:
@@
Before considering these metrics however, one requires a method to filter out ‘unacceptable’ samples. This is achieved using an admission function, $A$, which maps the outputs to $\{0,1\}$, effectively discarding samples that do not meet certain criteria. The choice of admission function can be tailored to specific tasks while still maintaining the coverage in Equation 1. Indeed, in Quach et al. (2024) we are presented with a range of instances of $A$. For instance, for the summarisation task $\mathrm{ROUGE}_{L}$ Lin (2004) scores were used, which focuses on measuring the longest common subsequence between two samples. This is used to compare sample summaries $y$ to the references $X$ being summarised (concretely, $A(y)=1$ if $\mathrm{ROUGE}_{L}(X,y)>0.35$ and 00 otherwise).
@@
Finally, after testing these quality and diversity, we test whether the confidence of the set is sufficient: $\mathcal{F}(\mathcal{C}_{i}(X))\geq\lambda_{3}$. If this is true, we break the loop and return $\mathcal{C}_{i}(X)$ as the conformal set. If not, we repeat the loop for the next set $\mathcal{C}_{i+1}(X)$. Next, we present the calibration procedure of Quach et al. (2024) to explain how to find appropriate constraints $\lambda$.
@@ 3.1 Calibration
First of all one identifies a calibration set $\mathcal{D}_{\mathrm{cal}}$, often a subset of validation or training data. This data is used to calibrate the values of $\lambda$ to ensure one’s samples conform to a desired standard. The calibration procedure in Quach et al. (2024) is an adaptation of the Learn Then Test (LTT) procedure of Angelopoulos et al. (2022). The calibration procedure is basically a search over a space of weighted configurations $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3})$222 In general, there is a value of $\lambda$ for each metric, meaning that the configuration-tuple $\lambda$ can be of arbitrary length.. Given the set of all possible configurations $\Lambda=\{0,0.01,\ldots,0.99,1\}^{3}$, we consider for each $\lambda\in\Lambda$ and each possible risk-tolerance $\varepsilon\in\{0,0.01,0.02,\ldots,0.99,1\}$ one calculates the binomial tail bound $p$-value under the null-hypothesis $\mathcal{H}_{\lambda}:\mathbb{E}[L(\lambda)]>\varepsilon$. This gives us a $p$-value $p_{\lambda,\varepsilon}=p_{\lambda}$:
@@
Following this long sequence of calculations of $p$-values, one applies a family-wise error rate controlling algorithm to the set of $p$-values to return the set of valid parameters $\Lambda_{\mathrm{valid}}$. In Quach et al. (2024) this is implemented using the ‘Bonferroni correction’:
@@
where $S_{\lambda}(X_{i})$ is the total number of samples made before rejecting (e.g. 20), $S^{*}(X_{i})$ is the index of the first correct sample, and $[\cdot]^{+}=\mathrm{max}(\cdot,0)$. The numerator $[S_{\lambda}(X)-S^{*}(X)]^{+}$ quantifies how quickly the LLM generates a correct sample. Since samples are generated sequentially, we ideally want to minimise superfluous generations which is done by minimising this number. We take $\rho_{1},\rho_{2}$ to be $0.5$, as done by Quach et al. (2024). We refer to Quach et al. (2024) for more details on this algorithm.
@@ 4 Methodology
In our approach, CLM is employed to ASP programs sampled from the LLM such that the generated programs are not only syntactically valid but also, the answer sets are likely to contain the correct solution when processed by the clingo solver Gebser et al. (2018). We investigate how different types of calibration sets—ranging from those containing single-step reasoning examples to those with multiple-step reasoning examples—affect the performance and generalizability of CLM. This focus allows us to understand the importance of sample diversity in the calibration phase and its impact on the model’s ability to generalise across varied reasoning tasks.
Our core methodology involves translating natural language inputs from the StepGame dataset into ASP programs using the LLM. To facilitate the translation of natural language task descriptions into Answer Set Programming (ASP) code, we employ In-Context Learning (ICL) within our methodology. ICL exploits the inherent ability of large language models (LLMs) to discern patterns and structures from provided examples within the input prompt, enabling them to generate contextually appropriate outputs without explicit parameter fine-tuning. The entire ASP program is then sampled using CLM, producing a conformal set of programs. These ASP programs are then processed by clingo to derive answer sets, with the assumption that a correct program resides within the conformal set of ASP programs.
@@ 5.1 Dataset
For our experiments we generated a dataset of 465,975 datapoints split into 341,284 for training, 11,350 for validation and 113,341 for testing. Within each of these sets there is a distribution of different numbers of hops. The generation was performed using the asp-solution.py script333 https://github.com/Fangjun-Li/SpatialLM-StepGame/blob/main/asp_solution.py written for the data-generation in Li et al. (2024). Full details of the samples we used to ensure replicability will be provided in an online Github page upon acceptance.
@@ 5.3 Metrics and measures
For the second experiment, we implement an LLM-as-Judge metric, inspired by recent methodologies that leverage language models for evaluating and comparing generated outputs Liu et al. (2023); Zheng et al. (2023). This approach utilises a fine-tuned language model to assess the quality of generated ASP samples by computing the difference in average log probabilities between pairs of samples. Specifically, we fine-tuned the llama3.1-8B-Instruct model on approximately one-third of the training set. This subset includes a representative distribution of reasoning steps ranging from 1 to 24, ensuring that the model is exposed to diverse reasoning complexities during fine-tuning. The fine-tuning process was conducted using the LoRa adapter from the LlamaFactory repository Zheng et al. (2024), allowing efficient adaptation of the model with minimal computational overhead. Notably, the fine-tuned model achieves 100% accuracy on the task, ensuring reliable evaluation performance.
@@ 5.4 Calibration
To evaluate the impact of calibration set composition on the generalizability of Conformal Language Modelling (CLM), we employed two distinct calibration sets. The first calibration set comprises StepGame entries where the number of reasoning steps ranges from 1 to 5, encompassing a diverse array of reasoning complexities. The second calibration set is restricted to entries with a only 1 reasoning step. Both calibration sets were extracted from the validation subset and consist of 500 entries each, ensuring a balanced and sufficient sample size for reliable calibration.

### v2

@@ Abstract
In this paper, we examine the use of Conformal Language Modelling (CLM) alongside Answer Set Programming (ASP) to enhance the performance of standard open-weight LLMs on complex multi-step reasoning tasks. Using the StepGame dataset, which requires spatial reasoning, we apply CLM to generate sets of ASP programs from an LLM, providing statistical guarantees on the correctness of the outputs. Experimental results show that CLM significantly outperforms baseline models that use standard sampling methods, achieving substantial accuracy improvements across different levels of reasoning complexity. Additionally, the LLM-as-Judge metric enhances CLM’s performance, especially in assessing structurally and logically correct ASP outputs. However, calibrating CLM with diverse calibration sets did not improve generalisability for tasks requiring many more reasoning steps, indicating limitations in handling more complex tasks.  [^1] We thank Microsoft Research - Accelerating Foundation Models Research program, for the provision of Azure resources to access OpenAI models. AGC also acknowledges partial support from the Economic and Social Research Council (ESRC) under grant ES/W003473/1.
@@ 1 Introduction
Building automated reasoning systems is one of the cornerstone goals of research in artificial intelligence, with wide-reaching applications in a variety of domains. LLMs have demonstrated remarkable capabilities in various natural language processing tasks Min et al. (2021), yet exhibit critical limitations in a variety of reasoning-related tasks, particularly in multistep and spatially complex scenarios Yang et al. (2023); Li et al. (2024). Although these models process and generate natural language with unprecedented fluency, their reasoning capabilities are still limited by several key constraints.
@@
Most LLMs frequently struggle with verifiable, robust and interpretable reasoning. Experimental studies have shown significant variability in LLM responses to identical reasoning related questions which highlights inherent instability Mirzaee & Kordjamshidi (2023). For instance, sampling from an LLM’s responses to reasoning-based queries can produce widely varying answers with significantly divergent reasoning traces. This lack of robustness undermines their reliability in critical decision-making contexts Ji et al. (2023). Moreover, tasks that require multiple sequential reasoning steps, such as complex spatial reasoning, often expose fundamental weaknesses in LLM’s systematic reasoning capabilities Shi et al. (2022).
@@
Recent research has increasingly focused on neural-symbolic approaches that combine the strengths of LLMs in learning from data and rigorous reasoning capabilities of symbolic systems Mao et al. (2019); Yu et al. (2024). Specifically, Answer Set Programming (ASP), a type of declarative logic programming, has emerged as a promising candidate for enhancing LLMs’ reasoning capabilities (Yang et al., 2023). Recent studies have explored integrating LLMs with ASP to transform natural language into logical representations, particularly for spatial reasoning tasks (Yang et al., 2023; Li et al., 2024).
@@
However, existing approaches face significant challenges in generating reliable intermediate scaffolds: the logical representations that bridge natural language inputs and systematic symbolic reasoning systems. The translation of natural language into formal logic programs introduces substantial uncertainty, as LLMs can produce inconsistent or semantically incorrect translations Li et al. (2024). This variability may become a significant weakness in this reasoning pipeline, potentially compromising the reliability of downstream reasoning tasks Mirzaee & Kordjamshidi (2023).
@@
In this paper, we present an approach that exploits the recently proposed Conformal Language Modelling (CLM) Quach et al. (2024) to systematically address these scaffold generation challenges. Conformal prediction (CP), the foundational framework for CLM, offers unique advantages in uncertainty quantification Vovk et al. (2005); Angelopoulos et al. (2022). CLM has been applied to natural language generation tasks like summarisation and question-answering Quach et al. (2024). Although recent advances in conformal prediction for language generation show promise, significant questions remain regarding the scalability, generalisability and robustness of these approaches across reasoning domains Campos et al. (2024). In this paper, we present a thorough analysis of these open challenges through an empirical lens. We assess how calibration in CP affects generalisability on multi-hop reasoning tasks and evaluate the efficacy of adapting metrics for formal language generation (ASP). Additionally, we identify common error types in the ASP scaffolded LLM based reasoning pipelines and examine how these propagate. Our findings demonstrate the reliability and performance of neuro-symbolic reasoning systems, contributing to the development of more robust and interpretable reasoning frameworks.
@@ 2 Background
In this section we introduce the main components used in our work: conformal prediction, conformal language modelling and answer set programming.
@@
Conformal Prediction is a statistical framework introduced by Vovk et al. (2005) that provides a method for constructing prediction regions with finite-sample guarantees under minimal assumptions. CP can be described as a sampling method which, under certain mild assumptions, guarantees that samples taken from any distribution using this method will be ‘correct’ with a controlled probability. This theory has been well-received in machine learning, where providing guarantees on the correctness of predictions helps improve performance, in particular since inference using neural networks involves sampling from theoretically unknown distributions Campos et al. (2024). CP has already proven its utility and efficacy in various domains, such as in computer vision tasks Hechtlinger et al. (2018) and drug discovery Cortés-Ciriano & Bender (2019).
@@
Conformal Language Modelling (CLM), developed in Quach et al. (2024), is a recent generalisation of CP which allows conformal sampling of language models. CLM proposes a conformal approach to generating sets of responses with language models, allowing for statistical guarantees in the vast and complex output space of natural language generation. The main contribution of CLM is an algorithm that can control a specific risk: the expected indicator loss that the generated set contains at least one acceptable response according to a predefined admission function—while simultaneously maximising other quantities of interest such as quality, diversity, and confidence. However, it remains unclear how CLM affects generalisability, particularly because it has been primarily tested on natural language generation tasks like summarisation, and report generation, where verifying the accuracy of models is not always straightforward Campos et al. (2024). Applying CLM to domains with rigorous notions of correctness, such as formal language generation, can provide better insights into its handling of generalisability and potential limitations.
@@
Answer set programming is a type of declarative programming which is commonly used in formal representations of knowledge, constraint satisfaction tasks, and formal common-sense reasoning. Answer set programs consist of finite sets of facts and rules that are used to generate answer sets under stable model semantics, which allows for efficient encoding of complex problems, notably in domains like graph search and combinatorial optimization Lifschitz (2019). ASP’s utility extends to various industrial applications due to its expressive power and computational efficiency, for instance, in e-tourism Ricca et al. (2010), high-performance computing scheduling Gamblin et al. (2022), configuration management, diagnosis, and planning Erdem et al. (2016). Further, ASP has been used to bolster spatial reasoning in LLMs through neural-symbolic integration Yang et al. (2023).
@@ 3 Formal introduction to CLM
At its core, CLM functions as a calibrated rejection sampling algorithm, allowing one to form conformal sets of samples, $\mathcal{C}(X)$, from an LLM, given a prompt $X$, such that the expected loss of $\mathcal{C}(X)$ is guaranteed to be within a chosen bound. Concretely, for a given loss function $L$ and error tolerance $\delta\in(0,1)$, CLM ensures there exists a risk tolerance $\varepsilon\in(0,1)$ such that for any prompt $X$, the conformal set $\mathcal{C}(X)$ satisfies the Equation 1 with probability at least $1-\delta$:
@@
Before considering these metrics however, one requires a method to filter out ‘unacceptable’ samples. This is achieved using an admission function, $A$, which maps samples to $\{0,1\}$, effectively discarding samples that do not meet certain criteria. The choice of admission function can be tailored to specific tasks while still maintaining the coverage from Equation 1. Indeed, in Quach et al. (2024) we are presented with a range of instances of $A$. For instance,for the summarisation task $\mathrm{ROUGE}_{L}$ Lin (2004) scores were used, which focuses on measuring the longest common subsequence between two samples. This is used to compare sample summaries $y$ to the references $X$ being summarised (concretely, $A(y)=1$ if $\mathrm{ROUGE}_{L}(X,y)>0.35$ and 00 otherwise).
@@
Finally, after testing quality and diversity, we test whether the confidence of the set is sufficient: $\mathcal{F}(\mathcal{C}_{i}(X))\geq\lambda_{3}$. If this is true, we break the loop and return $\mathcal{C}_{i}(X)$ as the conformal set. If not, we repeat the loop for the next set $\mathcal{C}_{i+1}(X)$. Next, we present the calibration procedure of Quach et al. (2024) to explain how to find appropriate constraints $\lambda$ in the first place.
@@ 3.1 Calibration
First of all one identifies a calibration set $\mathcal{D}_{\mathrm{cal}}$, often a subset of validation or training data. This data is used to calibrate the values of $\lambda$ to ensure one’s samples conform to a desired standard. The calibration procedure in Quach et al. (2024) is an adaptation of the Learn Then Test (LTT) procedure of Angelopoulos et al. (2022). The calibration procedure is basically a search over a space of weighted configurations $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3})$333 In general, there is a value of $\lambda$ for each metric, meaning that the configuration-tuple $\lambda$ can be of arbitrary length.. Given the set of all possible configurations $\Lambda=\{0,0.01,\ldots,0.99,1\}^{3}$, we consider for each $\lambda\in\Lambda$ and each possible risk-tolerance $\varepsilon\in\{0,0.01,0.02,\ldots,0.99,1\}$ one calculates the binomial tail bound $p$-value under the null-hypothesis $\mathcal{H}_{\lambda}:\mathbb{E}[L(\lambda)]>\varepsilon$. This gives us a $p$-value $p_{\lambda,\varepsilon}=p_{\lambda}$:
@@
Following this lengthy sequence of calculations of $p$-values, one applies a family-wise error rate controlling algorithm to the set of $p$-values to return the set of valid parameters $\Lambda_{\mathrm{valid}}$. In Quach et al. (2024) this is implemented using the ‘Bonferroni correction’:
@@
where $S_{\lambda}(X_{i})$ is the total number of samples made before rejecting (e.g. 20), $S^{*}(X_{i})$ is the index of the first correct sample, and $[\cdot]^{+}=\mathrm{max}(\cdot,0)$. The numerator $[S_{\lambda}(X)-S^{*}(X)]^{+}$ quantifies how quickly the LLM generates a correct sample. Since samples are generated sequentially, we ideally want to minimise superfluous generations which is done by minimising this number. We take $\rho_{1},\rho_{2}$ to be $0.5$, as done by Quach et al. (2024) where the reader can find further detail.
@@ 4 Methodology
In our approach, CLM is employed to ASP programs sampled from the LLM such that the generated programs are not only syntactically valid but also, the answer sets are likely to contain the correct solution when processed by the clingo ASP solver Gebser et al. (2018). We investigate how different types of calibration sets—ranging from those containing single-step reasoning examples to those with multiple-step reasoning examples—affect the performance and generalisability of CLM. This focus allows us to understand the importance of sample diversity in the calibration phase and its impact on the model’s ability to generalise across varied reasoning tasks.
Our core methodology involves translating natural language inputs from the StepGame dataset into ASP programs using the LLM. To facilitate the translation of natural language task descriptions into Answer Set Programming (ASP) code, we employ In-Context Learning (ICL). ICL exploits the inherent ability of large language models (LLMs) to discern patterns and structures from provided examples within the input prompt, enabling them to generate contextually appropriate outputs without explicit parameter fine-tuning. The entire ASP program is then sampled using CLM, producing a conformal set of programs. These ASP programs are then processed by clingo to derive answer sets.
@@ 5.1 Dataset
For our experiments we generated a dataset of 465,975 datapoints split into 341,284 for training, 11,350 for validation and 113,341 for testing. Within each of these sets there is a distribution of different numbers of hops. The generation was performed using the asp-solution.py script444 https://github.com/Fangjun-Li/SpatialLM-StepGame/blob/main/asp_solution.py written for the data-generation in Li et al. (2024).
@@ 5.3 Metrics and measures
For the second experiment, we implement an LLM-as-Judge metric, inspired by recent methodologies that leverage language models for evaluating and comparing generated outputs Liu et al. (2023); Zheng et al. (2023). This approach uses a fine-tuned language model to assess the quality of generated ASP samples by computing the difference in average log probabilities between pairs of samples. Specifically, we fine-tuned the llama3.1-8B-Instruct model on approximately one-third of the training set. This subset includes a representative distribution of reasoning steps ranging from 1 to 24, ensuring that the model is exposed to diverse reasoning complexities during fine-tuning. The fine-tuning process was conducted using the LoRa adapter from the LlamaFactory repository Zheng et al. (2024), allowing efficient adaptation of the model with minimal computational overhead. Notably, the fine-tuned model achieves 100% accuracy on the task, ensuring reliable evaluation performance.
@@ 5.4 Calibration
To evaluate the impact of calibration set composition on the generalisability of Conformal Language Modelling (CLM), we used two distinct calibration sets. The first calibration set comprises StepGame entries where the number of reasoning steps ranges from 1 to 5, encompassing a diverse array of reasoning complexities. The second calibration set is restricted to entries with a only 1 reasoning step. Both calibration sets were extracted from the validation subset and consist of 500 entries each, ensuring a balanced and sufficient sample size for reliable calibration.